# Emergence of Abstractions: Concept Encoding and Decoding Mechanism for In-Context Learning in Transformers

### v2

@@ 5 Discussion
5.1 Implications of Concept Encoding-Decoding Mechanism
Our proposed concept encoding-decoding mechanism has several implications in light of recent works on understanding the mechanics of ICL (Mittal et al., 2024) and activation-steering methods (Bürger et al., 2024; Panickssery et al., 2024; Marshall et al., 2024).
@@ 5.1 Implications of Concept Encoding-Decoding Mechanism
It is yet puzzling how to categorize the types of ICL tasks LLMs can and cannot solve (Qiu et al., 2023; Dziri et al., 2023). An intuitive explanation is that the model can effectively encode the concepts frequently seen during pretraining (Razeghi et al., 2022; Li et al., 2024). In our experiments, we observe patterns consistent with this conjecture, where AND and OR, the more common logical operators in language, were encoded more accurately. However, under our proposed two-stage mechanism, we show the bottleneck in ICL tasks can exist in both levels of concept inference and subsequent decoding algorithm. Therefore, even if the model already learned the algorithm for a NOR operator, if the model cannot clearly distinguish the latent concept from the inputs, it will fail, and vice versa. As our experiments suggest in Section 4.3, when the model is failing at concept encoding, a different prescription of early layer finetuning for representational alignment is more beneficial.
@@
Mittal et al. (2024) investigate whether explicitly modeling the latent variables in in-context learning (ICL) outperforms implicit learning through ordinary autoregressive training with transformer. They draw the counterintuitive conclusion that explicit modeling does not enhance performance, albeit not worse; the underlying reasons for which remain unclear. In our work, we explains this specific observation by analyzing the extent to which implicit modeling (standard transformers) captures the true latent variables. Our findings show that transformers can inherently encode these latent variables without explicit regularization. Therefore, we propose that the comparable performance between explicit and implicit models arises not because modeling the latent variables is unhelpful, but because both types of models effectively learn them.
Why may activation-based interventions work?
Our proposed encoding-decoding mechanism offers a lens to understand the broader impact of concept representations in LLMs beyond in-context learning. Recent works (Bürger et al., 2024; Panickssery et al., 2024; Marshall et al., 2024) observed that representations related to concepts such as semantics, truthfulness, or even vulnerabilities like jailbreaks often emerge naturally during pretraining, suggesting that models inherently encode such abstractions. Our framework also provides insight into why activation-based interventions, such as activation steering, can effectively modulate model behavior: by directly perturbing the encoded representations, these methods influence the downstream decoding algorithms that guide generation. Understanding how these representations and algorithms interact in contexts beyond ICL could open exciting avenues for future research, particularly in designing more robust, interpretable, and aligned language models.
5.2 Limitations
A limitation of our work is that the experimental setup used in this study does not encompass tasks that require multi-step reasoning (Clusmann et al., 2023; Zupan et al., 1999; Hosseini et al., 2024). Although we analyze the concept encoding-decoding mechanism with varying levels of complexity in Appendix D, further studies are essential to apply our findings and insights to the real-world. Another limitation stems from our proposed CD metric. Since we measure the separability from one concept to another, for the measure to be meaningful, the distribution of tasks on which CD is computed needs careful design to consist of semantically similar, confusing tasks.
6 Acknowledgements
We thank the members of the Improbable AI lab for the helpful discussions and feedback on the paper. We also thank Yoon Kim, Ekin Akyürek, Tamar Shaham, and Ziming Liu for constructive feedback on the experimental design and narrative. We are grateful to MIT Supercloud and the Lincoln Laboratory Supercomputing Center for providing HPC resources. The research was supported in part by NSF CSGrad4US Graduate Fellowship, Sloan Foundation and Schmidt Polymath Award. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government.
@@ Appendix A Investigating Predictability of ICL Task Performance in Large-Scale Pretraining
Figure 9: Test accuracy and CD scores of POS Tagging across OLMo-7B (Groeneveld et al., 2024) checkpoints, from 1000 to 500000. .
Appendix B Concept Encoding
In this section, we formally define the Concept Encoding and Concept Decoding.
Definition 2 (Concept Encoding)
Let $M$ be a transformer model, $\mathcal{Z}={z_{1},z_{2},\ldots,z_{n}}$ be a set of latent concepts, and $D$ be in-context examples with arbitrary length $K$. A concept encoding is an internal mapping $E:\mathcal{D}\rightarrow\mathbb{R}^{d_{emb}}$, where $\mathbb{R}^{d_{emb}}$ is the intermediate representation over the model’s $d$-dimensional embedding space.
Definition 3 (Concept Decoding)
Given a transformer model $M$ with concept encoding $E$, a concept decoding is a transformer’s behavior that there exists a simple function G that can recover the original latent concept and condition the algorithm:
@@ Appendix B Concept Encoding
Definition 4 (Decodability)
For a given decoder $G:\mathbb{R}^{d_{emb}}\rightarrow\mathcal{Z}$ and a specific latent variable $z$, we define the decodability measures as follows:
@@
    Accuracy:
    $A(G,z)=P(G(E(z,D))=z)$
@@
    Distribution Similarity:
    $S(G,z)=D_{f}(P(\hat{z})||P(z))$
Our study suggests that in transformers, the encoder $E$ maps distinct latent variables $z$ to separable representations. The model then applies different algorithms based on the inferred $\hat{z}$. This separability suggests that the transformer is inherently biased toward having a simple decoder $G$. In our study, use the kNN classifier for a decoder, accuracy and for score.
Appendix C Synthetic ICL Experiment
C.1 Theoretical Error Bounds in sparse linear regressions
@@ C.2 Experimental Details
We adapt the conventional linear regression setup from Garg et al. (2022); von Oswald et al. (2023) to create latent bases $\bm{B}$ that we can interpret far more easily than $W$. We study this setting with $D=16$ dimensional with up to $K=20$ in-context examples. Each $B_{i}$ has a rank of 4 and is orthogonal with each other. We independently sample $W$ and $x_{i}$ for each new input sequence from $N(0,\bm{I_{D}})$ the noise ${\epsilon}\sim N(0,0.01)$. We add the sparsity constraints to the linear regression task to introduce the latent concept of sparsity basis $B$ that is easily interpretable and analyzable in their representations. With the sparsity constraints, we construct the graphical model $B\rightarrow W\rightarrow Y\leftarrow X$. This construction allows us to visualize the representations of each of the bases (latent concepts in this graph) by aggregating the representations across a set of $W$ and $(X,Y)$ pairs.
@@
C.3 Additional results
@@
C.4 Additional Analysis on Section 3.3
CD Over Training.
We quantified the CD score for the synthetic experiments shown in Figure 2 , with the results presented in Figure 11 and Figure 12. The CD scores for Basis 1 effectively capture the separation of representations observed at (a). An increase in CD scores correlates with a corresponding drop in MSE, as seen in Figure 2, supporting our hypothesis that the CD score can serve as a predictor for the predictability of CD.
Figure 11: CD score of synthetic experiments in Figure 2 over training. (a), (b), (c) denote the same training points in Figure 2.
Figure 12: CD score across layers at epoch 10, 20, 100 from the synthetic experiment in Figure 2.
@@
Appendix D Increasing complexity in synthetic experiments
@@
(b) CD score across layers at epoch = 5,10,20,50
@@
D.1 Experiment - More Orthogonal Bases
@@
D.2 Experiment - Overlapping Bases
@@
(b) CD score across layers at epoch = 5,10,20,50
(c) CD score between nonoverlapping bases sets(btw basis 1,2,3,4 and 5,6,7,8) at epoch = 5,10,20,50
@@
Appendix E Natural ICL Experiments
@@
E.1 Mechanistic Intervention Study from Section 4.1
@@
E.2 Generalization with Different Model Families and Scales
@@
Figure 19: CD score vs ICL performance across Gemma-2 models (2B/9B/27B) and Llama-3.1-70B. The positive correlation between CD and ICL performance seen in Llama-3.1-8B generalizes across different models and scales. The grey dashed lines are linear lines of best fit. These results suggest that the accuracy of concept encoding is closely coupled with downstream ICL performance.
@@
E.3 Pairwise Concept Decodability Comparison
Figure 20: Pairwise CD scores for POS Tagging and arithmetic tasks at 4 shot. Pairwise CD scores identifies the clustered tasks
Appendix F Prompting Experiments
@@ Results.
Figure 22: CD score across layers for POS tagging and bitwise arithemetic in Llama-3.1-8B for the prompting experiments. We include the true labels of the latent concept (i.e. “Find the first noun in the sentence.”). We detail the experimental setup in Appendix F.

### v3

@@ 5 Discussion
Our insights on the origins and mechanism of task vectors have several implications in light of recent works on understanding the mechanics of ICL (Mittal et al., 2024) and activation-steering methods (Bürger et al., 2024; Panickssery et al., 2024; Marshall et al., 2024).
@@
It is yet puzzling how to categorize the types of ICL tasks LLMs can and cannot solve (Qiu et al., 2023; Dziri et al., 2023). An intuitive explanation is that the model can effectively encode the tasks frequently seen during pretraining (Razeghi et al., 2022; Li et al., 2024). In our experiments, we observe patterns consistent with this conjecture, where AND and OR, the more common logical operators in language, were encoded more accurately. However, under our proposed two-stage mechanism, we show the bottleneck in ICL tasks can exist in both levels of task inference and subsequent decoding algorithms. Therefore, even if the model already learned the algorithm for a task, if the model cannot clearly distinguish the latent concept from the inputs, it will fail, and vice versa.
@@
Mittal et al. (2024) investigate whether explicitly modeling the latent variables in ICL outperforms implicit learning through ordinary autoregressive training with transformer. They draw the counterintuitive conclusion that explicit modeling does not enhance performance, albeit not worse; the underlying reasons for which remain unclear. In our work, we explain this specific observation by analyzing the extent to which implicit modeling (standard transformers) captures the true latent variables. Our findings show that transformers can inherently encode these latent variables without explicit regularization. Therefore, we propose that the comparable performance between explicit and implicit models arises not because modeling the latent variables is unhelpful, but because both types of models effectively learn them.
Limitations.
A limitation of our work is that the experimental setup used in this study does not encompass tasks that require multi-step reasoning (Clusmann et al., 2023; Zupan et al., 1999; Hosseini et al., 2024). Although we analyze the task encoding-decoding mechanism with varying levels of complexity in Appendix E, further studies are essential to apply our findings and insights to the real-world. Another limitation stems from our proposed TD metric. Since we measure the separability from one task to the others, for the measure to be meaningful, the distribution of tasks on which TD is computed needs careful design to consist of semantically similar, confusing tasks.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Investigating Predictability of ICL Task Performance in Large-Scale Pretraining
(a) POS Tagging
(b) Bitwise Arithematic
Figure 6: Test accuracy and TD scores for POS tagging and bitwise arithmetic tasks across OLMo-7B (Groeneveld et al., 2024) checkpoints (1000-500000 steps), evaluated using 4-shot prompts on 100 test examples per task.
Appendix B Task Encoding
In this section, we formally define the Task Encoding and Decoding.
Definition B.1 (Task Encoding).
Let $M$ be a transformer model, $\mathcal{Z}={z_{1},z_{2},\ldots,z_{n}}$ be a set of latent tasks, and $D$ be in-context examples with arbitrary length $K$. A task encoding is an internal mapping $E:\mathcal{D}\rightarrow\mathbb{R}^{d_{emb}}$, where $\mathbb{R}^{d_{emb}}$ is the intermediate representation over the model’s $d_{e}mb$-dimensional embedding space.
Definition B.2 (Task Decoding).
Given a transformer model $M$ with concept encoding $E$, a task decoding is a transformer’s behavior that there exists a simple function G that can recover the original latent concept and condition the algorithm:
@@ Appendix B Task Encoding
Definition B.3 (Decodability).
For a given decoder $G:\mathbb{R}^{d_{emb}}\rightarrow\mathcal{Z}$ and a specific latent variable $z$, the decodability measures how accurately the correct latent variable is inferred from representations. Representation is distributed as $E(z,D)$, and the inferred latent variable is $\hat{z}\equiv G(E(z,D))$.
@@
    One-hot Accuracy: $A_{\text{1-hot}}(z)=E[\mathbf{1}\bigl{[}\hat{z}=z\bigr{]}]$
@@
    f-divergence: $A_{f}(z)=D_{f}(\hat{z}\parallel z)$, where $f$ is some f-divergence metric
In our study, we employ the one-hot accuracy metric with a kNN classifier to report task decodability.
Appendix C Activation Patching
[…]
Appendix D Synthetic ICL Experiment
D.1 Theoretical Error Bounds in sparse linear regressions
@@ D.2 Experimental Details
We adapt the conventional linear regression setup from (Garg et al., 2022; von Oswald et al., 2023) to create latent bases $\bm{B}$ that we can interpret far more easily than $W$. We study this setting with $D=16$ dimensional with up to $K=20$ in-context examples. Each $B_{i}$ has a rank of 4 and is orthogonal with each other. We independently sample $W$ and $x_{i}$ for each new input sequence from $N(0,\bm{I_{D}})$ the noise $\epsilon\sim N(0,0.01)$. We add the sparsity constraints to the linear regression task to introduce the latent concept of sparsity basis $B$ that is easily interpretable and analyzable in their representations. With the sparsity constraints, we construct the graphical model $B\rightarrow W\rightarrow Y\leftarrow X$. This construction allows us to visualize the representations of each of the bases (latent concepts in this graph) by aggregating the representations across a set of $W$ and $(X,Y)$ pairs.
@@
D.3 Additional results
Perturbation analysis to study the causal relation between task encoding and performance over the course of training.
@@ D.3 Additional results
D.4 Additional Analysis on Section 3.3
TD Over Training.
We quantified the TD score for the synthetic experiments shown in Figure 2 , with the results presented in Figure 9 and Figure 10. The TD scores for Basis 1 effectively capture the separation of representations observed at (a). An increase in TD scores correlates with a corresponding drop in MSE, as seen in Figure 2, supporting our hypothesis that the TD score can serve as a predictor for the predictability of TD.
Figure 9: TD score of synthetic experiments in Figure 2 over training. (a), (b), (c) denote the same training points in Figure 2.
Figure 10: TD score across layers at epoch 10, 20, 100 from the synthetic experiment in Figure 2.
@@
Appendix E Increasing complexity in synthetic experiments
@@
(b) TD score across layers at epoch = 5,10,20,50
@@
E.1 Experiment - More Orthogonal Bases
@@
E.2 Experiment - Overlapping Bases
@@
(b) TD score across layers at epoch = 5,10,20,50
(c) TD score between nonoverlapping bases sets(btw basis 1,2,3,4 and 5,6,7,8) at epoch = 5,10,20,50
@@
Appendix F Mixture of Different Regression Families Experiment
To further validate the robustness of the concept encoding-decoding framework under more diverse and complex conditions, we conduct experiments with a mixture of regression families—linear, polynomial (degree 3), and sinusoidal regressions. Each regression type represents a distinct latent concept with fundamentally different functional relationships between inputs and outputs.
Experimental Setup
We construct datasets with equal proportions of each regression type. Input dimensionality remains consistent with previous setups ($D=16$), and training parameters follow the synthetic experiment protocol outlined in Appendix D.2.
Results
Results shown in Figure 16 indicate clear separability of latent representations corresponding to each regression family after training. Furthermore, we observe significant performance improvements when latent concepts are clearly encoded, confirming that models effectively infer and leverage concept-specific algorithms. This suggests that transformers naturally learn to adaptively switch between qualitatively distinct prediction algorithms based on their internal representations.
(a) Loss curves for different regression families.
(b) UMAP visualization of representations at convergence.
(c) TD scores across layers for each regression family.
Figure 16: Mixture of Different Regression Families Analysis. (a) Loss curves highlight performance improvements when latent concepts are clearly encoded. (b) UMAP visualization demonstrates distinct clusters corresponding to each regression family, indicating successful latent concept encoding. (c) TD scores across layers validate the emergence of clear representations for each regression type.
Verification of Head-Pruning Results
We validate our claim that distinct latent bases form separate representations by analyzing attention-head pruning effects in synthetic and mixture regression tasks. Attention-head pruning was conducted individually based on magnitude-based importance scores, removing each head one at a time.
Figure 18 shows that attention-head pruning in synthetic sparse linear regression tasks. Changes in mean squared error (MSE) were measured across 100 random sequences per basis (same setup as main manuscript Figure 2). Attention Importance Estimation (AIE) quantifies performance degradation. Layer 4 distinctly maps attention heads to specific bases (e.g., l4h5 for base 0 and l4h3 for base 2), supporting distinct algorithm implementations.
Figure 17 shows that attention-head pruning on mixture regression families. MSE changes assessed across 100 random sequences for each regression type (linear, leaky ReLU, sparse linear, quadratic). Attention heads are notably shared across linear, leaky ReLU, and sparse linear regressions (excluding quadratic), indicating structural similarity. This aligns with the previously observed ”common structure” (Kim et al.), confirming structural indistinguishability among these regression variants.
Overall, This confirms that distinct algorithms require dedicated representational capacities, whereas shared algorithms benefit from structural overlap.
Figure 17: Head-Pruning: Mixture Tasks. Minimal impact of individual attention-head pruning indicates robustness from shared algorithmic structures.
Figure 18: Head-Pruning: Synthetic Tasks. Individual attention-head pruning significantly affects synthetic task performance, demonstrating distinct basis sensitivity.
Appendix G Natural ICL Experiments
@@
G.1 TD score by Layers from Section 4.1
We present the TD scores across the 32 layers of LLaMA 3.1 8B for POS tagging and bitwise arithmetic tasks. We found that TD scores peak at layer 15 for POS tagging and layer 13 for bitwise arithmetic tasks, and we used these layers for measuring TD throughout Section 4. The observation that TD scores peak in the middle layers is consistent with the findings of (Hendel et al., 2023) and (Todd et al., 2023).
(a) TD Scores By Layers
Figure 19: TD Scores by layers and number of demonstrations. (a) Mean TD scores across layers for POS tagging and Bitwise arithmetic with 4-shot in-context examples, showing peak decodability in intermediate layers. (b) For POS tagging and (c) for Bitwise arithmetic, TD scores all increase with the number of demonstrations, but the improvement in TD noticeably varies by task.
G.2 Mechanistic Intervention Study from Section 4.1
@@
G.3 Generalization with Different Model Families and Scales
@@
Figure 21: TD score vs ICL performance across Gemma-2 models (2B/9B/27B) and Llama-3.1-70B. The positive correlation between TD and ICL performance seen in Llama-3.1-8B generalizes across different models and scales. The grey dashed lines are linear lines of best fit. These results suggest that the accuracy of concept encoding is closely coupled with downstream ICL performance.
@@
G.4 Extension to Recurrent Neural Network Architectures
To examine whether our task encoding-decoding framework generalizes beyond transformer-based models, we conduct experiments using a recurrent neural network (RNN) architecture. Specifically, we utilize a two-layer LSTM model with 512 hidden units per layer. The training and evaluation setups replicate those used for transformer experiments in Figure 4. Figure 22 shows that the LSTM-based architecture also demonstrates a positive correlation between task decodability (TD score) and in-context learning performance. Although absolute performance metrics differ from transformer-based models, the fundamental relationship between representation clarity and task performance remains consistent. These findings suggest that the task encoding-decoding mechanism is not exclusive to transformers, but also is applicable to sequential neural models more broadly.
Figure 22: TD scores vs ICL perofrmance for Mamba-2 8B model. The positive relationship between TD scores and accuracy in the Mamba-2 8B model suggests that the capacity of TD scores to serve as a proxy for task encoding-decoding processes might generalize to RNN-based language models.
G.5 Pairwise Concept Decodability Comparison
Figure 23: Pairwise TD scores for POS Tagging and arithmetic tasks at 4 shot. Pairwise TD scores identifies the clustered tasks
Appendix H Finetuning Experiments
We visualize the ICL test accuracy changes before and after finetuning the first and last 10 layers of the model on each of the tasks in Figure 24. These results confirm the hypothesis that, contrary to the common practice of finetuning the last layers for classification tasks for instance, finetuning the earlier layers directly improves the task encoding and thus the ICL task performance more than finetuning the latter layers.
(a) POS Tagging
(b) Bitwise Arithmetic
Appendix I Prompting Experiments
@@ Results.
Figure 25: TD score across layers for POS tagging and bitwise arithemetic in Llama-3.1-8B for the prompting experiments. We include the true labels of the latent concept (i.e. “Find the first noun in the sentence.”). We detail the experimental setup in Appendix I.